# BRACTIVE: A Brain Activation Approach to Human Visual Brain Learning

### v2

@@ Abstract
The human brain is a highly efficient processing unit, and understanding how it works can inspire new algorithms and architectures in machine learning. In this work, we introduce a novel framework named Brain Activation Network (BRACTIVE), a transformer-based approach to studying the human visual brain. The main objective of BRACTIVE is to align the visual features of subjects with corresponding brain representations via fMRI signals. It allows us to identify the brain’s Regions of Interest (ROI) of the subjects. Unlike previous brain research methods, which can only identify ROIs for one subject at a time and are limited by the number of subjects, BRACTIVE automatically extends this identification to multiple subjects and ROIs. Our experiments demonstrate that BRACTIVE effectively identifies person-specific regions of interest, such as face and body-selective areas, aligning with neuroscience findings and indicating potential applicability to various object categories. More importantly, we found that leveraging human visual brain activity to guide deep neural networks enhances performance across various benchmarks. It encourages the potential of BRACTIVE in both neuroscience and machine intelligence studies.
@@ 1 Introduction
Artificial intelligence (AI) algorithms [1, 2, 3, 4] have recently made significant capabilities one step forward to a human-like level, thanks to the development of deep learning. Inspired by human brain mechanisms, the feedforward and deep neural networks have greatly enhanced performance across various domains such as computer vision, natural language processing, speech recognition, etc [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Recent studies show that these methods are more likely to mirror the mechanism of the human brain. For instance, a relationship has been found between brain representations in the visual pathway and the hierarchical layers within Deep Neural Networks (DNNs) [17, 18, 19, 20, 21]. They have lighted a research direction involving both cognitive neuroscientists and AI researchers to uncover the mechanisms and complexities of the human brain and then potentially adapting these findings to AI models [22, 23, 24, 25].
@@
In one direction, brain-inspired representations offer a promising avenue for enhancing DNNs training. For example, incorporating brain data, such as Electroencephalogram (EEG) signals, has been shown to improve performance in classification and salient detection tasks [26]. Another study showed that representations simulating the primate primary visual cortex in DNN models significantly enhance robustness to adversarial attacks and image corruptions [27]. This approach leverages the inherent strengths of neural processing to guide and refine artificial models. Conversely, AI models possess unique capabilities that can significantly advance our understanding of neuroscience [28, 29, 30]. These models can simulate complex neural interactions and offer insights that are challenging to obtain through traditional experimental methods. Specifically, our research highlights the advantages of integrating deep learning techniques with neuroscience to elucidate the topological representations of objects within the brain. By mapping these representations, we can gain a deeper understanding of how the brain organizes and processes visual information.
@@
Contribution of This Work. In summary, the contributions in this paper are four-folds. First, we present a novel transformer-based Brain Activation Network (BRACTIVE) for human visual brain understanding. Apart from the prior work, our approach can identify the brain’s Regions of Interest (ROI) of multiple subjects automatically as in Fig. 1. Second, we propose a novel module named Subject of Interest Proposal (SOIP) to predict the interested subjects that participants are focusing on during the visual tasks. Third, since the labels for the proposed tasks are not available, we introduce the Subject of Interest Retrieval (SOIR) module to establish features of these subjects in visual and fMRI modalities. Finally, the Weighted SOI Loss function is presented to align the features of subjects across modalities. The experiments show ROIs for humans identified by BRACTIVE are closely aligned with the findings (floc-faces, floc-bodies) from neuroscience. It promises to bring valuable references for neuroscientists in studying visual brain functions. The code will be released.
@@ 2 Related Work
Decoding human brain representation has been one of the most popular research topics for a decade. In particular, cognitive neuroscience has made substantial advances in understanding neural representations originating in the primary Visual Cortex (V1) [31]. Indeed, the primary visual context is responded to process information related to oriented edges and colors. The V1 forwards the information to other neural regions, focusing more on complex shapes and features. These regions are largely overlapped with receptive fields such as V4 [32], before converging on object and category representations in the inferior temporal (IT) cortex [33]. Neuroimaging techniques, including Functional Magnetic Resonance Imaging(fMRI), Magnetoencephalography(MEG), and electroencephalogram(EEG), have been crucial in these studies. However, to replicate human-level neural representations that fully capture our visual processes, it is crucial to precisely monitor the activity of every neuron in the brain simultaneously. Consequently, recent efforts in brain representation decoding have focused on exploring the correlation between neural activity data and computational models. In this research direction, several studies [34, 35, 36, 37, 38] were presented to decode brain information. Recently, with the help of deep learning, the authors in [39, 40, 41, 42, 43, 44] presented the methods to reconstruct what humans see from fMRI signal using diffusion models. The authors in [45, 39] also explored patterns of fMRI signals. However, they often fall short of demonstrating or explaining the nature of these patterns.
@@
In addition, there are recent studies on the intersection between AI and cognitive neuroscience [46, 47, 48, 49, 43, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 46, 60, 13, 61, 62, 63, 64, 65]. Deep learning methods [66, 67, 68, 69, 70] have been applied to predict the neural responses. Some other studies inspired by biological mechanisms such as memory [71], coding theory, and attention [72, 73] are increasingly adopted in the AI field. Some recent methods have used neural activity data to guide the training of models [26, 74, 22, 75, 76]. They utilized EEG and fMRI signals to constrain the neural network to behave like the neural response in the visual cortex.
@@
Discussion: Neuroscience and computer science mutually benefit each other. Insights from the human cognitive system can lead to innovative AI algorithms, while advanced deep neural networks can enhance our exploration of the human brain. By delving deeper into the intricacies of how the brain processes information, learns, and makes decisions, researchers are paving the way for a new era of AI. This understanding can be translated into more efficient AI solutions that require less training data and can learn and adapt in ways that mirror human learning. According to neuroscience findings, V1 is responsible for edge, and orientation processing, while V4 can handle larger receptive fields. The object-specific ROIs such as floc-faces and floc-bodies are responsible for recognizing human subjects. This prompts the question: Which Regions of Interest (ROIs) are related to processing information of other subjects, such as animals, bicycles, cars, etc?. Therefore, this paper focuses on understanding how the human visual brain processes visual stimuli. Apart from the prior studies, our approach is not limited by predefined visual ROI from the neuroscience field. Instead, the proposed method can automatically explore more ROI for a specific subject.
@@ 3 The Proposed Brain Activation Network Approach
In this section, we detail the proposed method. In particular, we first briefly introduce vision encoder, fMRI encoder, and text encoder in Subsections 3.1, 3.2, 3.3, respectively. Then, the details of the Subject of Interest Proposal, Subject of Interest Retrieval modules, and overall objective loss functions are presented in Subsections 3.4, 3.5, 3.6, correspondingly.
Figure 2: The illustration of BRACTIVE and training objectives. The Subject of Interest Proposal proposes the interested subjects that the participants are focusing on. The Subject of Interest Retrieval is used to construct the corresponding features of the proposed subjects on visual and fMRI modalities. The Weighted SOI Loss is used to align the features of these subjects.
@@ 3.1 Vision Encoder
Given a visual stimulus $x\in\mathbb{R}^{h\times w\times c}$, we split $x$ into multiple non-overlap patches $\mathcal{P}=\{p_{i}\}_{i=0}^{N_{p}-1}$ where $N_{p}=|\mathcal{P}|=(h\times w)/{p^{2}}$ is the number of patches, and $p$ is the patch size. Let $\texttt{Enc}_{v}$ be a visual encoder, i.e., ViT, which receives $\mathcal{P}$ as the input. The vectors of latent and patches are represented as in the Eq. (1).
@@
In dealing with the high-dimensional fMRI 1D signal $f\in\mathbb{R}^{N_{F}}$, where $N_{F}$ might reach 100K, a direct strategy involves employing linear or fully connected layers [40] to model the signal and derive its latent embedding. However, this approach encounters two main limitations. First, fully connected layers struggle with high-dimensional features, making it difficult to extract meaningful information while burdening the model’s memory usage. Second, the current representation form of the fMRI, as a signal, makes it difficult to explore the local patterns of nearby voxels (Please refer to the Appendix/Supplementary sections for the details). To address these challenges, we project these signals into a flattened 2D space by leveraging the transformation function of fsaverage surface. Then we have $f_{2d}=\texttt{T}_{2d}(f)\in\mathbb{R}^{h_{2d}\times w_{2d}}$ where $f_{2d}$ is the flattened 2D form of the fMRI signal $f$, $h_{2d}$ and $w_{2d}$ are the height and width, correspondingly. Consequently, we apply a transformer-based encoder, denoted $\texttt{Enc}_{f}$, to extract its feature. The output features of the fMRI signals are represented as in Eq. (3).
@@
Figure 3: Detailed illustration of Subject of Interest Proposal (SOIP) (a) Subject of Interest Retrieval (SOIR) (b) module. It is similar to applying the SOIR module for the fMRI modality by replacing vision tokens $\textbf{P}_{i}$ by $\textbf{F}_{i}$, respectively
@@ 3.3 Text Encoder
Since $\mathbf{P}_{s_{k}}$ and $\mathbf{F}_{s_{k}}$ represent the visual and fMRI modalities of the $s_{k}$ subject, respectively, we need semantic labels in visual to construct these features. However, we need prior access to this information, posing a significant obstacle. To address this challenge, we propose utilizing the text modality to represent the $s_{k}$ for the following reasons. First, unlike the visual and fMRI modalities that require multiple patches or pixels and sub-signals/voxels to represent the subject, a single word is sufficient to describe the subject (e.g., cat, dog, etc.) in the text modality. Secondly, the features of the text are more discriminative than images, thanks to BERT-based architecture [bert]. Thus, text poses a strong representation of a subject.
@@ 3.4 Subject of Interest Proposal
1:  Input: The visual stimulus $x$, fMRI signal $f$, text description $t$. Let $k$ be the number of the subject of interest proposal. 2:  Output: The objective loss. $\triangleright$ Extract visual features 3:  $\mathcal{P}\leftarrow\texttt{Patchify}(x)$ 4:  $\mathbf{P}_{\texttt{CLS}},\left[\mathbf{P}_{0},\dots\mathbf{P}_{i},\dots\mathbf{P}_{N_{p}-1}\right]\leftarrow\texttt{Enc}_{v}(\mathcal{P})$triangleright Extract fMRI features 5:  $f_{2d}\leftarrow\texttt{T}_{2d}(f)$ 6:  $\mathbf{F}_{\texttt{CLS}},\left[\mathbf{F}_{0},\dots\mathbf{F}_{i},\dots\mathbf{F}_{N_{r}-1}\right]\leftarrow\texttt{Enc}_{f}(f_{2d})$triangleright Extract text features 7:  $\mathbf{T}_{\texttt{CLS}},\left[\mathbf{T}_{0},\dots\mathbf{T}_{i},\dots\mathbf{T}_{N_{c}-1}\right]\leftarrow\texttt{Enc}_{t}(t)$triangleright Subject of interest proposal 8:  $\mathbf{T}_{\texttt{SOI}},G\leftarrow\texttt{SOIP}(\mathbf{T}_{\texttt{CLS}},k)$triangleright Subject of interest retrieval 9:  $\mathbf{P}_{\texttt{SOI}}\leftarrow\texttt{SOIR}(\mathbf{T}_{\texttt{SOI}},\mathbf{P}_{v})$ 10:  $\mathbf{T}_{\texttt{SOI}}\leftarrow\texttt{SOIR}(\mathbf{F}_{\texttt{SOI}},\mathbf{F}_{f})$triangleright Global contextual loss 11:  $\mathcal{L}_{g}\leftarrow\texttt{TripContras}(\mathbf{T}_{\texttt{CLS}},\mathbf{P}_{\texttt{CLS}},\mathbf{F}_{\texttt{CLS}})$triangleright Subject of interest loss 12:  $\mathcal{L}_{m}\leftarrow 0$ 13:  for $k^{\prime}=1,\cdots,k$ do 14:     $\mathbf{T}_{s_{k}},\mathbf{P}_{s_{k}},\mathbf{F}_{s_{k}}\leftarrow\mathbf{T}_{\texttt{SOI}}\left[k^{\prime}\right],\mathbf{P}_{\texttt{SOI}}\left[k^{\prime}\right],\mathbf{F}_{\texttt{SOI}}\left[k^{\prime}\right]$ 15:     $\mathcal{L}_{m}\leftarrow\mathcal{L}_{m}+G\left[k^{\prime}\right]*\texttt{WSOILoss}(\mathbf{T}_{s_{k}},\mathbf{P}_{s_{k}},\mathbf{F}_{s_{k}})$ 16:  end for 17:  $\mathcal{L}=\lambda_{g}\mathcal{L}_{g}+\lambda_{m}\mathcal{L}_{m}$ 18:  return $\mathcal{L}$
@@
Let $\mathcal{M}_{s}$ be a set of subjects within a visual stimulus that the participant is focusing. It is crucial to recognize that $\mathcal{M}_{s}$ varies for each individual. For instance, in a stimulus containing both dogs and cats, a person fond of dogs will focus solely on the regions containing dogs, while cat lovers will concentrate on the cats’ regions. Consequently, identifying $\mathcal{M}_{s}$ poses a challenge since we need to possess this predefined set of subjects of interest in advance. To address this challenge, we introduce the SOIP module to identify $\mathcal{M}_{s}$.
@@
Where $I$ is a vector stating the location of the token in the $t$ and $G$ is a vector representing the corresponding probability. $\mathbf{T}_{\texttt{SOI}}$ is the a feature set of $\mathcal{M}_{s}$ where each element is the token feature $\textbf{T}_{i}$ as defined in the Eq. (5) where $i\in I$.

### v3

@@ Abstract
The human brain is a highly efficient processing unit, and understanding how it works can inspire new algorithms and architectures in machine learning. In this work, we introduce a novel framework named Brain Activation Network (BRACTIVE), a transformer-based approach to studying the human visual brain. The primary objective of BRACTIVE is to align the visual features of subjects with their corresponding brain representations using functional Magnetic Resonance Imaging (fMRI) signals. It enables us to identify the brain’s Regions of Interest (ROIs) in the subjects. Unlike previous brain research methods, which can only identify ROIs for one subject at a time and are limited by the number of subjects, BRACTIVE automatically extends this identification to multiple subjects and ROIs. Our experiments demonstrate that BRACTIVE effectively identifies person-specific regions of interest, such as face and body-selective areas, aligning with neuroscience findings and indicating potential applicability to various object categories. More importantly, we found that leveraging human visual brain activity to guide deep neural networks enhances performance across various benchmarks. It encourages the potential of BRACTIVE in both neuroscience and machine intelligence studies.
@@ 1 Introduction
Artificial intelligence (AI) algorithms [1, 2, 3, 4] have recently advanced significantly, bringing capabilities one step closer to a human-like level, thanks to the development of deep learning. Inspired by human brain mechanisms, the feedforward and deep neural networks have greatly enhanced performance across various domains such as computer vision, natural language processing, speech recognition, etc [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Recent studies show that these methods are more likely to mirror the mechanism of the human brain. For instance, a relationship has been found between brain representations in the visual pathway and the hierarchical layers within Deep Neural Networks (DNNs) [17, 18, 19, 20, 21]. They have initiated a research direction involving both cognitive neuroscientists and AI researchers to uncover the mechanisms and complexities of the human brain, with the potential to adapt these findings to AI models [22, 23, 24, 25].
@@
In one direction, brain-inspired representations offer a promising avenue for enhancing DNN training. For example, incorporating brain data, such as Electroencephalogram (EEG) signals, has been shown to improve performance in classification and salient detection tasks [26]. Another study showed that representations simulating the primate primary visual cortex in DNN models significantly enhance robustness to adversarial attacks and image corruptions [27]. This approach leverages the inherent strengths of neural processing to guide and refine artificial models. Conversely, AI models possess unique capabilities that can significantly advance our understanding of neuroscience [28, 29, 30]. These models can simulate complex neural interactions, offering insights that are challenging to obtain through traditional experimental methods. Specifically, our research highlights the advantages of integrating deep learning techniques with neuroscience to elucidate the topological representations of objects within the brain. By mapping these representations, we can gain a deeper understanding of how the brain organizes and processes visual information.
@@
Contributions of This Work. In summary, the contributions in this paper are four-folds. First, we present a novel transformer-based Brain Activation Network (BRACTIVE) [^1] The code of this work will be released. for understanding the human visual brain. Apart from the prior work, our approach can automatically identify the brain’s Regions of Interest (ROIs) for multiple subjects, as shown in Fig. 1. Second, we propose a novel module, named Subject of Interest Proposal (SOIP), to predict the subjects that participants focus on during visual tasks. Third, since the labels for the proposed tasks are not available, we introduce the Subject of Interest Retrieval (SOIR) module to establish features of these subjects in visual and fMRI modalities. Finally, the Weighted SOI Loss function is presented to align the features of subjects across modalities. The experiments show ROIs for humans identified by BRACTIVE are closely aligned with the findings (floc-faces, floc-bodies) from neuroscience. It promises to bring valuable references for neuroscientists in studying visual brain functions.
@@ 2 Related Work
Decoding human brain representation has been one of the most popular research topics for a decade. In particular, cognitive neuroscience has made substantial advances in understanding neural representations originating in the primary Visual Cortex (V1) [31]. Indeed, the primary visual context responds to the processing of information related to oriented edges and colors. The V1 forwards the information to other neural regions, focusing more on complex shapes and features. These regions are overlapped mainly with receptive fields such as V4 [32], before converging on object and category representations in the inferior temporal (IT) cortex [33]. Neuroimaging techniques, including Functional Magnetic Resonance Imaging (fMRI), Magnetoencephalography (MEG), and electroencephalogram (EEG), have been crucial in these studies. However, to replicate human-level neural representations that fully capture our visual processes, it is crucial to precisely monitor the activity of every neuron in the brain simultaneously. Consequently, recent efforts in brain representation decoding have focused on exploring the correlation between neural activity data and computational models. In this research direction, several studies [34, 35, 36, 37, 38] were presented to decode brain information. Recently, with the help of deep learning, authors in [39, 40, 41, 42, 43, 44] have presented methods to reconstruct what humans see from fMRI signals using diffusion models. The authors in [45, 39] also explored patterns of fMRI signals. However, they often fall short of demonstrating or explaining the nature of these patterns.
@@
In addition, there are recent studies on the intersection between AI and cognitive neuroscience [46, 47, 48, 49, 43, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 46, 71, 13, 72, 73, 74, 75, 76]. Deep learning methods [77, 78, 79, 80, 81] have been applied to predict the neural responses. Other studies, inspired by biological mechanisms such as memory [82], coding theory, and attention [83, 84], are increasingly being adopted in the AI field. Some recent methods have used neural activity data to guide the training of models [26, 85, 22, 86, 87]. They utilized EEG and fMRI signals to constrain the neural network to behave in a manner similar to the neural response in the visual cortex.
@@
Discussion: Neuroscience and computer science mutually benefit each other. Insights from the human cognitive system can lead to innovative AI algorithms, while advanced deep neural networks can enhance our exploration of the human brain. By delving deeper into the intricacies of how the brain processes information, learns, and makes decisions, researchers are paving the way for a new era of AI. This understanding can be translated into more efficient AI solutions that require less training data and can learn and adapt in ways that mirror human learning. According to neuroscience findings, V1 is responsible for edge and orientation processing, while V4 can handle larger receptive fields. The object-specific ROIs such as floc-faces and floc-bodies are responsible for recognizing human subjects. This prompts the question: Which Regions of Interest (ROIs) are related to processing information of other subjects, such as animals, bicycles, cars, etc?. Therefore, this paper focuses on understanding how the human visual brain processes visual stimuli. Apart from the prior studies, our approach is not limited by predefined visual ROI from the neuroscience field. Instead, the proposed method can automatically explore more ROI for a specific subject.
@@ 3 The Proposed Brain Activation Network Approach
In this section, we detail the proposed method as illustrated in Fig. 2. In particular, we first briefly introduce the vision encoder, fMRI encoder, and text encoder in Subsections 3.1, 3.2, 3.3, respectively. Then, the details of the Subject of Interest Proposal, Subject of Interest Retrieval modules, and overall objective loss functions are presented in Subsections 3.4, 3.5, 3.6, correspondingly.
Figure 2: The illustration of BRACTIVE and training objectives. The Subject of Interest Proposal identifies the interested subjects that participants are focusing on. The Subject of Interest Retrieval is used to construct the corresponding features of the proposed subjects on visual and fMRI modalities. The Weighted SOI Loss is used to align the features of these subjects.
@@ 3.1 Vision Encoder
Given a visual stimulus $x\in\mathbb{R}^{h\times w\times c}$, we split $x$ into multiple non-overlap patches $\mathcal{P}=\{p_{i}\}_{i=0}^{N_{p}-1}$ where $N_{p}=|\mathcal{P}|=(h\times w)/{p^{2}}$ is the number of patches, and $p$ is the patch size. Let $\texttt{Enc}_{v}$ be a visual encoder, i.e., ViT, which receives $\mathcal{P}$ as the input. The vectors of latent and patches are represented as in Eq. (1).
@@
In dealing with the high-dimensional fMRI 1D signal $f\in\mathbb{R}^{N_{F}}$, where $N_{F}$ might reach 100K, a direct strategy involves employing linear or fully connected layers [40] to model the signal and derive its latent embedding. However, this approach encounters two main limitations. First, fully connected layers struggle with high-dimensional features, making it difficult to extract meaningful information while burdening the model’s memory usage. Second, the current representation form of the fMRI, as a signal, makes it difficult to explore the local patterns of nearby voxels (Please refer to the Appendix/Supplementary for the details). To address these challenges, we project these signals into a flattened 2D space by leveraging the transformation function of fsaverage surface. Then we have $f_{2d}=\texttt{T}_{2d}(f)\in\mathbb{R}^{h_{2d}\times w_{2d}}$ where $f_{2d}$ is the flattened 2D form of the fMRI signal $f$, $h_{2d}$ and $w_{2d}$ are the height and width, correspondingly. Consequently, we apply a transformer-based encoder, denoted $\texttt{Enc}_{f}$, to extract its feature. The output features of the fMRI signals are represented as in Eq. (3).
@@
Figure 3: Detailed illustration of (a) Subject of Interest Proposal (SOIP) and (b) Subject of Interest Retrieval (SOIR) module. It is similar to applying the SOIR module for the fMRI modality by replacing vision tokens $\textbf{P}_{i}$ by $\textbf{F}_{i}$, respectively.
@@ 3.3 Text Encoder
Since $\mathbf{P}_{s_{k}}$ and $\mathbf{F}_{s_{k}}$ represent the visual and fMRI modalities of the $s_{k}$ subject, respectively, we need semantic labels in visual to construct these features. However, we require prior access to this information, which poses a significant obstacle. To address this challenge, we propose utilizing the text modality to represent the $s_{k}$ for the following reasons. First, unlike the visual and fMRI modalities that require multiple patches or pixels and sub-signals/voxels to represent the subject, a single word is sufficient to describe the subject (e.g., cat, dog, etc.) in the text modality. Secondly, the features of the text are more discriminative than those of images, thanks to the BERT-based architecture [1]. Thus, text poses a strong representation of a subject.
@@ 3.4 Subject of Interest Proposal
1:  Input: The visual stimulus $x$, fMRI signal $f$, text description $t$. Let $k$ be the number of the subject of interest proposal. 2:  Output: The objective loss. $\triangleright$ Extract visual features 3:  $\mathcal{P}\leftarrow\texttt{Patchify}(x)$ 4:  $\mathbf{P}_{\texttt{CLS}},\left[\mathbf{P}_{0},\dots\mathbf{P}_{i},\dots\mathbf{P}_{N_{p}-1}\right]\leftarrow\texttt{Enc}_{v}(\mathcal{P})$ $\triangleright$ Extract fMRI features 5:  $f_{2d}\leftarrow\texttt{T}_{2d}(f)$ 6:  $\mathbf{F}_{\texttt{CLS}},\left[\mathbf{F}_{0},\dots\mathbf{F}_{i},\dots\mathbf{F}_{N_{r}-1}\right]\leftarrow\texttt{Enc}_{f}(f_{2d})$ $\triangleright$ Extract text features 7:  $\mathbf{T}_{\texttt{CLS}},\left[\mathbf{T}_{0},\dots\mathbf{T}_{i},\dots\mathbf{T}_{N_{c}-1}\right]\leftarrow\texttt{Enc}_{t}(t)$ $\triangleright$ Subject of interest proposal 8:  $\mathbf{T}_{\texttt{SOI}},G\leftarrow\texttt{SOIP}(\mathbf{T}_{\texttt{CLS}},k)$ $\triangleright$ Subject of interest retrieval 9:  $\mathbf{P}_{\texttt{SOI}}\leftarrow\texttt{SOIR}(\mathbf{T}_{\texttt{SOI}},\mathbf{P}_{v})$ 10:  $\mathbf{T}_{\texttt{SOI}}\leftarrow\texttt{SOIR}(\mathbf{F}_{\texttt{SOI}},\mathbf{F}_{f})$ $\triangleright$ Global contextual loss 11:  $\mathcal{L}_{g}\leftarrow\texttt{TripContras}(\mathbf{T}_{\texttt{CLS}},\mathbf{P}_{\texttt{CLS}},\mathbf{F}_{\texttt{CLS}})$ $\triangleright$ Subject of interest loss 12:  $\mathcal{L}_{m}\leftarrow 0$ 13:  for $k^{\prime}=1,\cdots,k$ do 14:   $\mathbf{T}_{s_{k}},\mathbf{P}_{s_{k}},\mathbf{F}_{s_{k}}\leftarrow\mathbf{T}_{\texttt{SOI}}\left[k^{\prime}\right],\mathbf{P}_{\texttt{SOI}}\left[k^{\prime}\right],\mathbf{F}_{\texttt{SOI}}\left[k^{\prime}\right]$ 15:   $\mathcal{L}_{m}\leftarrow\mathcal{L}_{m}+G\left[k^{\prime}\right]*\texttt{WSOILoss}(\mathbf{T}_{s_{k}},\mathbf{P}_{s_{k}},\mathbf{F}_{s_{k}})$ 16:  end for 17:  $\mathcal{L}=\lambda_{g}\mathcal{L}_{g}+\lambda_{m}\mathcal{L}_{m}$ 18:  return $\mathcal{L}$
@@
Let $\mathcal{M}_{s}$ be a set of subjects within a visual stimulus that the participant is focusing on. It is crucial to recognize that $\mathcal{M}_{s}$ varies for each individual. For instance, in a stimulus containing both dogs and cats, a person fond of dogs will focus solely on the regions containing dogs, while cat lovers will concentrate on the cats’ regions. Consequently, identifying $\mathcal{M}_{s}$ poses a challenge since we need to possess this predefined set of subjects of interest in advance. To address this challenge, we introduce the SOIP module to identify $\mathcal{M}_{s}$.
@@
Where $I$ is a vector stating the location of the token in the $t$ and $G$ is a vector representing the corresponding probability. $\mathbf{T}_{\texttt{SOI}}$ is the a feature set of $\mathcal{M}_{s}$ where each element is the token feature $\textbf{T}_{i}$ as defined in the Eq. (5) where $i\in I$. The visualization of the SOIP module is illustrated in Fig. 3.